# Deep Transfer Learning: Model Framework and Error Analysis

### v1

@@ Abstract
This paper presents a framework for deep transfer learning, which aims to leverage information from multi-domain upstream data with a large number of samples ($n$) to a single-domain downstream task with a considerably smaller number of samples ($m$), where $m\ll n$, in order to enhance performance on downstream task. Our framework has several intriguing features. First, it allows the existence of both shared and specific features among multi-domain data and provides a framework for automatic identification, achieving precise transfer and utilization of information. Second, our model framework explicitly indicates the upstream features that contribute to downstream tasks, establishing a relationship between upstream domains and downstream tasks, thereby enhancing interpretability. Error analysis demonstrates that the transfer under our framework can significantly improve the convergence rate for learning Lipschitz functions in downstream supervised tasks, reducing it from $\tilde{O}(m^{-\frac{1}{2(d+2)}}+n^{-\frac{1}{2(d+2)}})$ (“no transfer”) to $\tilde{O}(m^{-\frac{1}{2(d^{*}+3)}}+n^{-\frac{1}{2(d+2)}})$ (“partial transfer”), and even to $\tilde{O}(m^{-1/2}+n^{-\frac{1}{2(d+2)}})$ (“complete transfer”), where $d^{*}\ll d$ and $d$ is the dimension of the observed data. Our theoretical findings are substantiated by empirical experiments conducted on image classification datasets, along with a regression dataset.
@@ 1 Introduction
Training a model solely on one or multiple analogous tasks can impede its performance when applied to a distinct yet similar task [1, 2], due to the presence of data bias and task bias. By recognizing that multiple diverse tasks may share a common representation, transfer learning emerges as a crucial solution. The relevant literature on transfer learning [3, 4] provides valuable insights into its applications and discussions. However, the existing literature lacks systematic research on the following issues, particularly within the framework of deep learning:
@@
    Given that the representations of different datasets should not be identical and there should be shared as well as specific features, how can we design a model framework to accurately characterize this?
@@
In this paper, we aim to address the aforementioned questions by proposing a deep learning framework for transfer learning. The goal is to derive feature representations from multi-domain upstream data, characterized by a large sample size, and apply them to a single-domain downstream task with a considerably smaller number of samples, in order to improve performance on the downstream task. Our model accomplishes this in three key steps. Firstly, it permits the presence of common and distinct features across domains in the upstream data and automatically identifies them. Secondly, it explicitly expresses which upstream features contribute to the downstream tasks, establishing connection among domains and enhancing interpretability. Moreover, a $d$-dimensional nonlinear function $Q^{*}$ is introduced for the downstream task to learn features that are independent of the upstream representation, determining the level of difficulty in transfer learning. Particularly, in the case where $Q^{*}=0$, the upstream representation is sufficient to the downstream task, transfer from upstream to downstream is facile, leading to a parametric convergence rate. On the other hand, when specific information is encoded in the non-zero $Q^{*}$, the transfer problem becomes more challenging. Nevertheless, with the aid of the representation learned from the upstream data, our task is simplified to the estimation of an independent $Q^{*}$ of upstream representation, resulting in a convergence rate that surpasses the pace of the Empirical Risk Minimization (ERM) estimator for the downstream task. Importantly, we emphasize that our model framework seamlessly adapts to these scenarios, accommodating varying degrees of information transfer and yielding distinct convergence rates for the downstream task.
The main contributions in this paper are as follows:
  • We introduce an adaptive deep transfer learning framework. Initially, we acquire deep representations from the upstream dataset through identifiable models, ensuring sufficiency and invariance. In instances of “complete transfer”, where the acquired representations offer sufficient information for downstream prediction, we only need to learn an $r$-dimensional vector function for the downstream task. In case that “partial transfer”, an additional learning for a low dimensional function is required.
  • Our error analysis reveals that the convergence speed, when learning a $\beta$-smooth function with over-parameterized neural networks for the downstream task, can manifest at distinct rates: slow $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$, moderate $\mathcal{O}(m^{-\frac{\beta}{2(d^{*}+1+2\beta)}})$ with $d^{*}\ll d$ or fast $\mathcal{O}(m^{-1/2})$. These rates correspond to “no transfer” that the downstream task necessitates learning a $d$-dimensional function, “partial transfer” problems where the downstream task entails learning an $d^{*}$-dimensional function, and “complete transfer” that the downstream task involves solving a sparse-penalized least squares problem, respectively.
1.1 Related Work
Extensive research has been conducted on single-source transfer learning under covariate shift and posterior shift, encompassing both parametric and nonparametric models. Theoretical investigations into transfer learning within parametric models, such as high-dimensional linear regression [5] or generalized linear models [6], have been thoroughly explored. Nonparametric transfer learning in regression with posterior shift has been addressed by [7, 8], as well as in the work of [9]. This area has also been explored by [10, 11] in the scenario of covariate shift. Nonparametric transfer learning in classification, has been studied in general by [12, 13, 14]. Theoretical results and adaptive procedures have been developed for both the covariate shift setting [15, 16] and the posterior drift setting [17, 18]. Minimax rates are derived under posterior drift by [8] and under covariate shift by [19], respectively. However, it is noteworthy that even in the extensive theoretical literature mentioned above, deep transfer learning, even for the single-source (a special case of our study with $p=1$), has not been thoroughly explored. Non-parametric analysis of transfer learning involving shared representations across diverse tasks have been explored, see [20, 21, 22, 23] and the reference therein. In downstream task, task-specific linear functions and nonlinear functions have been studied in [20, 22] and [21, 23], respectively. In these studies, their theories consistently demonstrate the improvement of transfer learning in terms of sampling efficiency under the assumption that sufficient representations of downstream task can be obtained from the upstream domains. In addition, these above mentioned theory is developed for well-specified neural network models without considering the approximation error as we did in this paper.
The concept of a shared invariant representation across domains is not only applicable in transfer learning but also extends to various related domains and tasks, including domain adaptation [24, 25, 26], domain generalization [27, 28, 29, 30], and meta-learning [31]. Recently, [32] provided theoretical insights into domain-invariant representations within the context of multiple-source domain adaptation and domain generalization. In this paper, we learned deep representations from the upstream dataset through identifiable models, ensuring not only invariance but also sufficiency. Another point sets our method apart from the aforementioned ones is that we allow for the existence of not only shared representations but also specific features, making our model more flexible and practical. Additionally, the loss function we used to train this representation differs significantly from the above ones.
Learning a pre-trained model with huge data set and then fine-tuning on specific task [33, 34, 35, 36, 37, 38, 39, 40, 41, 42] has empirically proven to be a highly effective strategy in recent years. Our framework for multi-domain transfer learning can also be interpreted along these lines, wherein learning representations from multiple domains can be viewed as pre-training, and refining downstream predictions can be seen as fine-tuning. Consequently, the theoretical results established in this paper also contribute to enhancing our understanding of pre-training and fine-tuning processes.
@@ 1.1 Related Work
The remainder of this paper is structured as follows: Preliminary details about the framework, including essential definitions and model setup, are provided in Section 1. Our main theoretical results are presented in Section 3. Experimental results are showcased in Section 4, followed by a summary in Section 5. Technical proofs are provided in the appendix.
@@ 2 Preliminaries
Table 1: Summary of notations
$\mathcal{D}=\{(\mathbf{X}_{i},Y_{i},S_{i})\}_{i=1}^{n}$ dataset for upstream
$\mathbb{P}(\mathbf{x},y,\mathbf{s})$ the unknown distribution of upstream data
$(\mathbf{X},Y,S)$ r.v.s (random variables) for upstream data
$\mathcal{X}$,$\mathcal{Y}$,$\mathcal{S}$ support of upstream data $\mathbf{X}$, $Y$, $\mathbf{s}$ respectively
$\mathcal{T}=\{(\mathbf{X}_{T,i},Y_{T,i})\}_{i=1}^{m}$ dataset for downstream
$\mathbb{P}_{0}$ the unknown distribution of downstream data
($\mathbf{X}_{T},Y_{T}$) the pair of r.v.s (random variables) $(X_{T},Y_{T})$ with its joint distribution $\mathbb{P}_{0}$
$\varepsilon_{1},\varepsilon_{2}$ bounded random noises: $|\varepsilon_{i}|\leq 1,i\in[2]$
$\mathcal{U}_{r}$ a random vector with each coordinate being mutually independent and
have uniform distribution $U[0,1]$
$\{\xi_{i}\}_{i=1}^{n}$ the i.i.d random samples drawn from $\mathcal{U}_{r}$
$n$ the number of upstream data
$m$ the number of downstream data
$p$ the number of different domain
$n_{s}$ the number of upstream data from domain $s$
$d$ the dimension of input data $\mathbf{X}$
$d^{*}$ the reduced dimension of downstream input data $\mathbf{X}_{T}$
$r$ the dimension of intrinsic representation
($\mathbf{X},Y$) the combine r.v.s(random variables) pairs with
@@
$\mathbf{X}_{S}$, $Y_{S}$ the r.v.s(random variables) from domain $S$
$\mathbf{X}\perp\!\!\!\perp Y$ for r.v.s (random variables) $\mathbf{X}$ and $Y$, $\mathbf{X}$ is independent with $Y$
$\mathbf{X}\perp\!\!\!\perp Y\mid Z$ for r.v.s (random variables) $\mathbf{X}$, $Y$ and $Z$, condition on $Z$,
$\mathbf{X}$ is independent with $Y$
$\mathbf{h}^{*}_{s}(\mathbf{X})$ the oracle representation for domain $s$ with $\mathbf{X}_{s}\perp\!\!\!\perp Y_{s}\mid\mathbf{h}^{*}_{s}(\mathbf{X})$
$\mathbf{h}^{*}(\mathbf{X})$ the oracle representation satisfied $\mathbf{X}\perp\!\!\!\perp Y\mid\mathbf{h}^{*}(\mathbf{X})$ and $S\perp\!\!\!\perp\mathbf{h}^{*}(\mathbf{X})$
$\mathbf{F}^{*}_{S}$ the oracle sparse transformation vector for domain $S$
$\mathbf{F}_{T}^{*}$ the oracle sparse transformation vector for downstream data
$Q^{*}$ an additional $\mathbb{R}^{d}\to\mathbb{R}$ function to be learned at downstream task
and satisfy $Q^{*}(X_{T})\perp\!\!\!\perp\mathbf{h}^{*}(X_{T})$
$\mathbf{A}^{*}$ the oracle dimensionality reduction matrix for downstream input data
to control downstream task complexity
$q^{*}$ the oracle $\mathbb{R}^{d^{*}}\mapsto\mathbb{R}$ function satisfy $Q^{*}(X_{T})=q^{*}(\mathbf{A}^{*}X_{T})$
$\mathbf{F}\in\mathbb{R}^{pr}$ the shorthand for $(\mathbf{F}_{1},\ldots,\mathbf{F}_{p})$
[…]
$\|\cdot\|_{1}$ a vector’s $\ell_{1}$ norm or the sum of the $\ell_{1}$ norms of the columns of a matrix
$[q]$ shorthand for integer sets $\{1,\ldots,q\}$
$\|f\|_{L}$ the Lipschitz constant of function $f$
$\mathcal{H}$ a neural network(NN) function class of features mapping $\mathbb{R}^{d}\to\mathbb{R}^{r}$
$\mathcal{G}_{1}$,$\mathcal{H}_{1}$ a NN function class of lipschitz-constrained mapping $\mathbb{R}^{r}\to\mathbb{R}$
$\mathcal{Q}$ a NN function class during tuning mapping $\mathbb{R}^{d}\to\mathbb{R}$
$\lambda,\tau,\mu,\kappa,\chi,\zeta$ regularization parameters
$\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$ a measure used to evaluate independence between $\mathbf{h}(\mathbf{X})$ and $S$,
in this work we use the distance covariance
$\widehat{\mathcal{V}}_{n}[\mathbf{h}(\mathbf{X}),S]$ the empirical form of $\mathcal{V}[\mathbf{h}(\mathbf{X}),S]$
$\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})$ the 1-Wasserstein distance used to match the distribution of $\mathbf{h}(\mathbf{X})$ and
$\mathcal{U}_{r}$, in this work we use its dual form to calculate it
$\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})$ the empirical dual form of $\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})$ by replacing the
searching function class by norm-constrained NN
$R(\mathbf{F},\mathbf{h})$ the population risk of learned function $(\mathbf{F},\mathbf{h})$;
i.e, $\mathbb{E}[\|Y_{S}-\mathbf{F}_{S}\mathbf{h}(\mathbf{X}_{S})\|^{2}]+\lambda\mathcal{V}[\mathbf{h}(\mathbf{X}),S]+\tau\mathcal{W}_{1}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})+\mu\|\mathbf{F}\|_{1}$
$\hat{R}(\mathbf{F},\mathbf{h})$ the empirical version of $R(\mathbf{F},\mathbf{h})$;
i.e, $\sum_{i=1}^{p}\sum_{j=1}^{n_{p}}\frac{1}{n}(Y_{i,j}-\mathbf{F}_{i}\mathbf{h}(X_{i,j}))^{2}+\lambda\widehat{\mathcal{V}}_{n}[\mathbf{h}(\mathbf{X}),S]+$
$\tau\hat{\mathcal{W}_{1}}(\mathbf{h}(\mathbf{X}),\mathcal{U}_{r})+\mu\|\mathbf{F}\|_{1}$
$(\hat{\mathbf{F}},\hat{\mathbf{h}})$ the estimated sparse vector $\hat{\mathbf{F}}=(\hat{\mathbf{F}}_{1},\ldots,\hat{\mathbf{F}}_{p})$ and estimated feature
representation $\hat{\mathbf{h}}$; i.e, $(\hat{\mathbf{F}},\hat{\mathbf{h}})=\mathop{\rm argmin}_{(\mathbf{F},\mathbf{h})\in\mathcal{B}_{pr}(B)\circ\mathcal{H}}\hat{R}(\mathbf{F},\mathbf{h})$
$\ell(\cdot,\cdot)$ the general loss for regression or classification
[…]
^$\hat{\mathcal{L}}(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F}_{T})$ the empirical risk of learned $(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F}_{T})$;
i.e, $\sum_{i=1}^{m}\ell(\mathbf{F}_{T}\hat{\mathbf{h}}(\mathbf{X}_{T,i})+q(\mathbf{A}\mathbf{X}_{T,i}),Y_{T,i})/m+\kappa\hat{\mathcal{V}}_{m}[\hat{\mathbf{h}},Q_{q,\mathbf{A}}]+$
   $\chi\|\mathbf{F}_{T}\|_{1}+{\zeta\|\mathbf{A}\|_{F}^{2}}$
$\mathcal{L}(\mathbf{h},Q_{q,\mathbf{A}},\mathbf{F}_{T})$ the population risk of learned $(\hat{\mathbf{h}},Q_{q,\mathbf{A}},\mathbf{F}_{T})$;
i.e, $\mathcal{L}(\mathbf{h},Q_{q,\mathbf{A}},\mathbf{F}_{T})=\mathbb{E}_{X_{T},Y_{T}}[\hat{\mathcal{L}}(\mathbf{h}(X_{T}),q(\mathbf{A}X_{T}),\mathbf{F}_{T})]$
$\tilde{O}(\cdot)$ Asymptotic notation used to hide polylogarithmic factors in all relevant
parameters of the problem
@@ 2.1.1 Setup
where $\mathbf{X}\perp\!\!\!\perp Y\mid\mathbf{h}^{*}(\mathbf{X})$ indicates that the information of $\mathbf{X}$ to predict $Y$ is sufficiently encoded in $\mathbf{h}^{*}(\mathbf{X})$, and $S\perp\!\!\!\perp\mathbf{h}^{*}(\mathbf{X})$ signifies the need for the representation to be invariant across all domains in the upstream dataset. Under (1), we use the following regression model for identifying and estimating $\mathbf{h}^{*}(\mathbf{X})$:
@@
where the sparse vector $\mathbf{F}^{*}_{S}\in\mathbb{R}^{1\times r}$ represents the transformation applied to the representation $\mathbf{h}^{*}(\mathbf{X}_{S})$ for domain $S$, $\varepsilon_{1}$ is a bounded noise with mean zero. The sparsity of $\mathbf{F}^{*}_{S}$ is used to identify the feature subset $\mathbf{h}^{*}_{S}(\mathbf{X}_{S})(\subset\mathbf{h}^{*}(\mathbf{X}_{S}))$ that capture the relationship between $\mathbf{X}_{S}$ and $Y_{S}$. The sparsity sets our framework apart from existing transfer learning methods by allowing features to vary or be shared across domains, rather than exclusively being shared. With the sparsity of $\mathbf{F}_{S}$, model (3) also provides a more parsimonious and interpretable representation for each domain. Domains with different feature sets enable us to better explore their interdomain differences. For the downstream data, we focus on single-target domain transfer learning. Let the data set $\mathcal{T}=\{(\mathbf{X}_{T,i},Y_{T,i})\}_{i=1}^{m}$ consist of $m$ samples of the random variable pair $(X{T},Y_{T})$, which are i.i.d. from an unknown distribution $\mathbb{P}_{0}$, which is also supported over $\mathcal{X}\times\mathcal{Y}$. We assume the supervised downstream task data following the regression model:

### v3

@@ Abstract
This paper introduces a framework for deep transfer learning aimed at improving performance on a single-domain downstream task with a limited sample size ($m$) by leveraging information from multi-domain upstream data with a significantly larger sample size ($n$), where $m\ll n$. Our framework offers several intriguing features. First, it allows the existence of both shared and domain-specific features across multi-domain data and provides a framework for automatic identification, achieving precise transfer and utilization of information. Second, the framework explicitly identifies upstream features that contribute to downstream tasks, establishing clear relationships between upstream domains and downstream tasks, thereby enhancing interpretability. Error analysis shows that our framework can significantly improve the convergence rate for learning Lipschitz functions in downstream supervised tasks, reducing it from $\tilde{O}(m^{-\frac{1}{2(d+2)}}+n^{-\frac{1}{2(d+2)}})$ (for “no transfer”) to $\tilde{O}(m^{-\frac{1}{2(d^{*}+3)}}+n^{-\frac{1}{2(d+2)}})$ (for “partial transfer”), and even to $\tilde{O}(m^{-1/2}+n^{-\frac{1}{2(d+2)}})$ (for “complete transfer”), where $d^{*}\ll d$ and $d$ is the dimension of the observed data. Our theoretical findings are supported by empirical experiments on image classification and regression datasets.
Keywords: transfer learning, nonparametric analysis, convergence rate.
@@ 1 Introduction
Transfer learning (TL) is a powerful technique that improves model performance in target domains with limited data by leveraging knowledge from related source domains. This approach addresses the common challenge of obtaining large datasets necessary for high predictive accuracy. TL has been successfully applied in diverse fields, including natural language processing (Ruder et al., 2019), medical image analysis (Kora et al., 2022), drug discovery (Cai et al., 2020), and materials science (Jha et al., 2019; Kim et al., 2021; Ju et al., 2021).
Due to its notable successes in practical applications, extensive research has focused on theory of transfer learning under conditions of covariate shift and posterior shift, addressing both parametric and nonparametric models. Specifically, covariate shift refers to changes in the marginal distributions of covariates between the source and target domains, while posterior shift involves changes in the relationships between covariates and outcomes across these domains. Theoretical investigations into transfer learning within parametric models, such as high-dimensional linear regression (Li et al., 2022), generalized linear models (Bastani, 2021; Li, Zhang, Cai and Li, 2023; Tian and Feng, 2023) or gaussian graphical models with false discovery rate control (Li, Cai and Li, 2023), have been thoroughly explored. However, once parametric models are specified, they often lack the flexibility to handle the complex posterior shifts encountered in real-world settings, leading to insufficient information transfer due to the varying expressions in parameters for the different models.
Nonparametric transfer learning with covariate shift has been addressed by Huang et al. (2006); Wen et al. (2014); Wang et al. (2016); Cai and Pu (2024); Schmidt-Hieber and Zamolodtchikov (2024) for regression tasks and by Ben-David et al. (2006); Blitzer et al. (2007); Mansour et al. (2009) for classification tasks. Theoretical results have been developed for both the covariate shift setting (Shimodaira, 2000; Sugiyama et al., 2007) and the posterior drift setting (Cai and Wei, 2021; Reeve et al., 2021). Paritcularly, minimax rates are derived under posterior drift by Cai and Pu (2024) and under covariate shift by Kpotufe and Martinet (2021), respectively. These nonparametric methods are primarily based on traditional nonparametric approximation techniques, such as local polynomial methods (Fan and Gijbels, 2018), spline approximation (Schumaker, 2007), and reproducing kernel techniques (Berlinet and Thomas-Agnan, 2011; Lv et al., 2018). In practice, these methods often face the “curse of dimensionality” when the covariate dimension $d$ is moderate to high, for instance, when $d>3$.
Thanks to their powerful function-fitting capability, well-designed architectures, effective training algorithms, and high-performance computing technologies, nonparametric analysis based on deep neural networks (DNNs) has achieved tremendous success. Deep transfer learning, which uses DNNs to transfer learned representations across tasks, has been studied in the literature. For example, Du et al. (2020), Tripuraneni et al. (2020), Tripuraneni et al. (2021), Chen et al. (2021), and the references therein develop theories that consistently demonstrate the improvement of deep transfer learning in terms of sampling efficiency, under the assumption that sufficient representations of the downstream task can be obtained from the upstream domains and certain transferability conditions. However, these studies do not account for cases where downstream tasks may require more than representations which the upstream tasks can provide.
While prior research has advanced our understanding of transfer learning, the existing literature lacks systematic research on the following issues, particularly within the framework of deep learning:
@@
    Given that the representations of different datasets should not be identical and there should be shared as well as domain-specific features, how can we design a model framework to accurately estimate and characterize them?
@@
In this paper, we aim to address the aforementioned questions by proposing a deep learning framework for transfer learning. Our model accomplishes this in three key aspects. Firstly, our model permits the presence of common and distinct features across domains in the upstream and downstream data and provides framework to automatically identifies them. Secondly, it explicitly expresses which upstream features contribute to the downstream tasks, establishing connection among domains and enhancing interpretability. In addition, a non-linear function $Q^{*}$ is introduced for the downstream task to learn features that are independent of the upstream representation, determining the level of difficulty in transfer learning. Particularly, in the case where $Q^{*}=0$, the upstream representation is sufficient to the downstream task, transfer from upstream to downstream is facile, leading to “complete transfer” and a parametric convergence rate. On the other hand, when specific information is encoded in the non-zero $Q^{*}$, the transfer problem becomes more challenging. Our error analysis reveals that the convergence speed varies from slow $\mathcal{O}(m^{-\frac{\beta}{2(d+1+2\beta)}})$, to moderate $\mathcal{O}(m^{-\frac{\beta}{2(d^{*}+1+2\beta)}})$ with $d^{*}\ll d$. These rates correspond to “no transfer” that the downstream task necessitates learning a $d$-dimensional function, and “partial transfer” problems where the downstream task entails learning an $d^{*}$-dimensional function. Importantly, we emphasize that our model framework seamlessly adapts to these scenarios, accommodating varying degrees of information transfer and yielding distinct convergence rates for the downstream task.
@@
The remainder of this paper is structured as follows: Preliminary details about the framework, including essential definitions and model setup, are provided in Section 2. Our main theoretical results are presented in Section 3. Experimental results are showcased in Section 4, followed by a discussion in Section 5. Technical proofs are provided in the Supplementary Materials.
@@ 2.1.1 Setup
where $\mathbf{X}\perp\!\!\!\perp Y\mid\mathbf{h}^{*}(\mathbf{X})$ indicates that the information of $\mathbf{X}$ to predict $Y$ is sufficiently encoded in $\mathbf{h}^{*}(\mathbf{X})$, and $S\perp\!\!\!\perp\mathbf{h}^{*}(\mathbf{X})$ signifies the need for the representation to be invariant across all domains in the upstream dataset. Under (1), and employing the shorthand notation $\mathbf{X}_{S},Y_{S}$ to represent the triplet $(\mathbf{X},Y,S)$, we use the following regression model for identifying and estimating $\mathbf{h}^{*}(\mathbf{X})$:
@@
where the sparse vector $\mathbf{F}^{*}_{S}\in\mathbb{R}^{1\times r}$ represents the transformation applied to the representation $\mathbf{h}^{*}(\mathbf{X}_{S})$ for domain $S$, $\varepsilon_{1}$ is a bounded noise with mean zero. The sparsity of $\mathbf{F}^{*}_{S}$ is used to identify the feature subset $\mathbf{h}^{*}_{S}(\mathbf{X}_{S})(\subset\mathbf{h}^{*}(\mathbf{X}_{S}))$ that capture the relationship between $\mathbf{X}_{S}$ and $Y_{S}$. The sparsity sets our framework apart from existing transfer learning methods by allowing features to vary or be shared across domains, rather than exclusively being shared. With the sparsity of $\mathbf{F}_{S}$, model (3) also provides a more parsimonious and interpretable representation for each domain. Domains with different feature sets enable us to better explore their interdomain differences. For the downstream data, we focus on single-target domain transfer learning. Let the data set $\mathcal{T}=\{(\mathbf{X}_{T,i},Y_{T,i})\}_{i=1}^{m}$ consist of $m$ samples of the random variable pair $(X_{T},Y_{T})$, which are i.i.d. from an unknown distribution $\mathbb{P}_{0}$, which is also supported over $\mathcal{X}\times\mathcal{Y}$. We assume the supervised downstream task data following the regression model: